# Topological components of surface group representations and the Signature

### v1

@@ 4.2. Interior hyperbolic representations
By Lemmas 4.7 and 4.8, the restriction of $\phi$ to $\pi_{1}(M)$ satisfies the hypotheses, so there exists a path $\{\phi_{t}\}_{0\leq t\leq 1}$ in $\mathcal{R}_{hp}(\Sigma)$ such that $\phi_{1}(c)$ is hyperbolic and for each $c_{j}$, the elements $\phi_{t}(c_{j})$ remain conjugate to $\phi(c_{j})$ for all $t$. Let $\{U^{j}_{t}\}_{0\leq t\leq 1}$ be a path in $\mathrm{SL}(2,\mathbb{R})$ such that $\phi_{t}(c_{j})=U^{j}_{t}\phi(c_{j})(U^{j}_{t})^{-1}.$ We extend $\phi_{t}$ from $\pi_{1}(M)$ to $\pi_{1}(\Sigma)$ by defining $\phi_{t}(\gamma)=U^{j}_{t}\phi(\gamma)(U^{j}_{t})^{-1}$ for all $\gamma\in\pi_{1}(N_{j})$. This defines a continuous path $\{\phi_{t}\}_{0\leq t\leq 1}$ of representations in $\mathcal{R}_{hp}(\Sigma)$ such that $\phi_{1}(c)$ is hyperbolic. By a small perturbation, we may assume that $\phi_{1}\in\mathcal{R}^{\prime}_{hp}(\Sigma)$.
Now the restriction of $\phi_{1}$ to $\pi_{1}\left(\bigcup_{i>1}M_{i}\right)$ lies in $\mathcal{R}^{\prime}_{hp}(\Sigma^{\prime})$, where $\Sigma^{\prime}=\bigcup_{i>1}M_{i}$ is a subsurface of $\Sigma$. By the induction hypothesis, there exists a path in $\mathcal{R}_{hp}(\Sigma^{\prime})$ from this representation to one which maps each $c\in\partial M_{i}\setminus\partial\Sigma$ (for $i>1$) to a hyperbolic element. Furthermore, by the path-lifting property (see Lemma 4.9), this path can be extended to a path in $\mathcal{R}_{hp}(\Sigma)$. This completes the proof. ∎
4.3. The maximal Toledo invariant
@@
Let $\Sigma_{g,n}$ be a surface of genus $g$ with $n$ boundary components, and let $\mathrm{Par}(\Sigma_{g,n})$ denote the space of boundary parabolic representations. Then
@@
We first consider the case where $n\geq 2$. By Lemma 4.10, we may assume that $C_{0},D_{1},\dots,D_{n-2}$ are all hyperbolic.
@@
$\{\mathrm{T}^{-1}(|\chi(\Sigma_{g,n})|)\cap\mathrm{Par}(\Sigma_{g,n})\}=2^{2g+n-1}.$
@@
4.4. The case of $n$ is even
Let $\Sigma_{g,n}$ be a surface of genus $g$ with $n$ boundary components. Let $\mathrm{Hyp}(\Sigma_{g,n})$ denote the space of boundary hyperbolic representations in $\mathrm{Hom}(\pi_{1}(\Sigma_{g,n}),\mathrm{SL}(2,\mathbb{R}))$. For a hyperbolic element $C\in\mathrm{SL}(2,\mathbb{R})$, recall the sigma map is $\sigma(C)=\frac{1}{2}(1+\mathrm{sgn}(\mathrm{tr}(C)))$.
@@ 4.4. The case of $n$ is even
$\{\widehat{\mathcal{P}}_{g,n}\}=8^{n/2}\cdot\{\#\widehat{\mathcal{H}}_{g,n/2}\}=8^{n/2}\cdot 2^{n/2-1}(4g+n-3)=2^{2n-1}(4g+n-3).$
Note that $\widehat{\mathcal{P}}_{0,n}=\mathcal{P}_{0,n}$, which follows that
Corollary 4.17.
The number of connected components of $\mathrm{Par}(\Sigma_{0,n})$ is
$\{\mathrm{Par}(\Sigma_{0,n})\}=2^{2n-1}(n-3),$
where $n\geq 4$ is even.
@@
Let $\phi\in\mathrm{Hom}(\pi_{1}(\Sigma_{g,n}),\mathrm{SL}(2,\mathbb{R}))$ be a representation such that $\phi(c_{1}),\dots,\phi(c_{n-1})$ are parabolic and $\phi(c_{n})$ is hyperbolic. Then there exists a deformation $\phi(t)$ of $\phi$ such that the conjugacy classes of $\phi(t)(c_{1}),\dots,\phi(t)(c_{n-1})$ remain fixed, the trace $\mathrm{tr}(\phi(t)(c_{n}))$ varies monotonically, and $\phi(1)(c_{n})$ is parabolic with $\mathrm{tr}(\phi(1)(c_{n}))=2\mathrm{sgn}(\mathrm{tr}(\phi(c_{n})))$.
@@
We first assume $n\geq 2$. By a small perturbation fixing the conjugacy classes of $\phi(c_{1}),\dots,\phi(c_{n-1})$, we may assume that $\phi(\Pi_{i=1}^{g}[a_{i},b_{i}]c_{1}\cdots c_{n-2})$ is either elliptic or hyperbolic. Up to conjugation, we may write
@@
$\mathrm{T}(\phi)=\mathrm{T}(\psi)+\tfrac{1}{2}s_{a},$
@@
Let $\phi\in\mathrm{Par}(\Sigma_{g,n})$ with even $n\geq 2$. Then $\phi$ and $\gamma\phi$ lie in the same connected component of $\mathrm{Par}(\Sigma_{g,n})$ for any $\gamma\in\Lambda_{g,n}$ if and only if $\mathrm{T}(\psi)$ is neither maximal nor minimal.
@@
For any $\phi_{0},\phi_{1}\in\{\phi\}$, the proof of Lemma 4.15 shows that the corresponding representations $\psi_{0}$ and $\psi_{1}\in\mathrm{Hyp}(\Sigma_{n/2})$ have the same Toledo invariant and the same sigma values.
@@
If $\mathrm{T}(\psi_{0})=\mathrm{T}(\psi_{1})$ is maximal or minimal, then $\psi_{0}$ and $\gamma\cdot\psi_{1}$ are connected for some $\gamma\in\Lambda_{g,n}$. From the proof of Lemma 4.20, we know that each of the elements $\phi^{\prime}_{i}(a_{j})$ and $\phi^{\prime}_{i}(b_{j})$, for $i=1,2$ and $1\leq j\leq g$, cannot be elliptic. Since $\phi^{\prime}_{0}$ and $\phi^{\prime}_{1}$ are connected, the corresponding elements $\psi(a_{j})=\phi^{\prime}_{i}(a_{j})$ and $\psi(b_{j})=\phi^{\prime}_{i}(b_{j})$ must lie in the closure of the same hyperbolic region. This implies that $\gamma=\mathrm{id}$, and hence $\psi_{0}$ and $\psi_{1}$ are connected.
@@
Hence, the number of connected components of the space $\mathrm{Par}(\Sigma_{g,n})$ of boundary-parabolic representations, assuming $n\geq 2$ is even and $(g,n)\neq(0,2)$, is given by
@@ 4.5.1. Connected components of $\mathrm{HP}(\Sigma_{g,n})$
Let $a=(a_{1},a_{2},a_{3})\in\{0,1\}\times\{0,1\}\times\{\pm 1,\pm\sqrt{-1}\}$. Then the subset $\sigma^{-1}(a)\subset\mathrm{HP}(\Sigma_{0,3})$ is non-empty. Moreover, $\sigma^{-1}(a)$ is connected if $\sum_{i=1}^{3}\mathrm{Re}(a_{i})$ is odd, and has two connected components if $\sum_{i=1}^{3}\mathrm{Re}(a_{i})$ is even. In particular, the space $\mathrm{HP}(\Sigma_{0,3})$ has $24$ connected components.
@@
Suppose $c\neq 0$. If $a_{1}+a_{2}+\mathrm{Re}(a_{3})$ is even, then $\mathrm{tr}(C_{2}C_{3})=(-1)^{a_{1}}(a+d+cs_{3})$. One can choose parameters $a,d,c$ such that $a+d>2$ and $a+d+cs_{3}<-2$. Hence $C_{1}=(C_{2}C_{3})^{-1}$ is hyperbolic and satisfies $\sigma(C_{1})=a_{1}$. Similarly, when $a_{1}+a_{2}+\mathrm{Re}(a_{3})$ is odd, we can choose $a,d,c$ such that $a+d+cs_{3}>2$ and again $C_{1}$ is hyperbolic with the correct sigma value. In both cases, we conclude that $\sigma^{-1}(a)$ is non-empty.
@@
Now suppose that $\sum_{i=1}^{3}\mathrm{Re}(a_{i})$ is even. Let $\phi,\phi^{\prime}\in\sigma^{-1}(a)$ be two representations with corresponding parameters $c$ and $c^{\prime}$ such that $cc^{\prime}>0$. Then a deformation as above shows that $\phi$ and $\phi^{\prime}$ lie in the same connected component of $\sigma^{-1}(a)$. However, if $cc^{\prime}<0$, then any path $\phi(t)$ connecting $\phi$ and $\phi^{\prime}$ would necessarily pass through some $t_{0}$ with $c(t_{0})=0$. At this point, the trace $\mathrm{tr}(C_{1}(t_{0}))$ would correspond to a representation with $\sum_{i=1}^{3}\mathrm{Re}(a_{i})$ odd, which contradicts the requirement that $\phi(t)\in\sigma^{-1}(a)$ for all $t$. Thus, $\phi$ and $\phi^{\prime}$ lie in distinct connected components.
This completes the proof. ∎
@@
By the definition of the Toledo invariant, we have $\mathrm{T}(\phi)=0$ for any $\phi\in\sigma^{-1}(a)$ with $\sum_{i=1}^{3}\mathrm{Re}(a_{i})$ odd. When $\sum_{i=1}^{3}\mathrm{Re}(a_{i})$ is even, the set $\sigma^{-1}(a)$ has two connected components, corresponding respectively to $\mathrm{T}^{-1}(\pm 1)$. Therefore, if $\mathrm{HP}(\Sigma_{0,3})\cap\sigma^{-1}(a)\cap\mathrm{T}^{-1}(k)$ is non-empty, then it must be connected.
@@
In both cases, the conjugacy classes of $\phi(c_{i})$, $1\leq i\leq n-2$, are unchanged during the deformation, $\phi(c_{n-1})$ is perturbed slightly, and $\phi(c_{n})$ becomes hyperbolic with sigma value $|\mathrm{Re}(\sigma(\phi(c_{n})))|$ for all $t>0$. Hence the Toledo invariant remains constant throughout the deformation.
@@
is well-defined and is a double covering.
@@
We now prove that $\Psi$ is a double covering. Consider a pair of pants $(C_{1},C_{2},C_{3})$ with $C_{2}$ and $C_{3}$ hyperbolic. By (3.3), there is a deformation of $C_{3}$ to a parabolic element, keeping the conjugacy class of $C_{1}$ and the sigma values of $C_{2}$ fixed. During this deformation, the trace of $C_{3}$ maintains its sign, and the Toledo invariant remains constant. In particular, if $C_{1}$ and $C_{2}$ do not commute, then by (3.2), $C_{3}$ can be deformed into two distinct parabolic elements with sigma values $\pm\sqrt{\mathrm{sgn}(\mathrm{tr}(C_{3}))}$.
Hence, for any $\phi^{\prime}\in\mathrm{Hyp}(\Sigma_{g,n})$ with $n\geq 2$, up to a small perturbation, we may assume that $\phi^{\prime}|_{\pi_{1}(P_{n})}$ is non-abelian. Then there exist at least two representations $\phi_{0},\phi_{1}\in\mathrm{HP}(\Sigma_{g,n})$ with different sigma values on $c_{n}$ such that $\Psi(\{\phi_{0}\})=\Psi(\{\phi_{1}\})=\{\phi^{\prime}\}$. Thus, each fiber of $\Psi$ has at least two elements.
To show that each fiber has at most two elements, suppose $\phi_{0},\phi_{1}\in\mathrm{HP}(\Sigma_{g,n})\cap\sigma^{-1}(a)\cap\Psi^{-1}(\{\phi^{\prime}\})$. Then $\phi_{0}^{\prime}$ and $\phi_{1}^{\prime}$ are connected by a path $\phi^{\prime}(t)\in\mathrm{Hyp}(\Sigma_{g,n})$. Reversing the deformation from $\phi^{\prime}(t)(c_{n})$ back to a parabolic element with fixed sigma value (as in Lemma 4.25), we obtain a path $\phi(t)$ in $\mathrm{HP}(\Sigma_{g,n})\cap\sigma^{-1}(a)$ connecting $\phi_{0}$ and $\phi_{1}$. Thus, the fiber $\Psi^{-1}(\{\phi^{\prime}\})\cap\sigma^{-1}(a)$ is connected, and so has at most one connected component. Since sigma values on $c_{n}$ distinguish between $\phi_{0}$ and $\phi_{1}$, the entire fiber has exactly two elements.
The argument for the case $n=1$ is analogous: the deformation occurs on the boundary curve $c_{1}$, while the part corresponding to $\Pi_{i=1}^{g-1}[a_{i},b_{i}]$ remains unchanged, and only a slight perturbation is performed on $a_{g},b_{g}]$, which remains hyperbolic throughout the deformation. This allows the construction to proceed in the same way. This proves that $\Psi$ is a double covering. ∎
@@
(4.21) $\{\mathrm{HP}(\Sigma_{g,n})\}=2\cdot\#\{\mathrm{Hyp}(\Sigma_{g,n})\}=2^{n}(4g+2n-5)+4^{g}\cdot 2^{n+1},$
for any $(g,n)\neq(1,1)$.

### v3

@@ 4.2.1. Interior hyperbolic representations
By Lemmas 4.8 and 4.7, the restriction of $\phi$ to $\pi_{1}(M)$ satisfies the hypotheses, so there exists a path $\{\phi_{t}\}_{0\leq t\leq 1}$ in $\mathcal{R}^{\prime}_{h}(M)$ such that $\phi_{1}(c)$ is hyperbolic and for each $c_{j}$, the elements $\phi_{t}(c_{j})$ remain conjugate to $\phi(c_{j})$ for all $t$. Let $\{U^{j}_{t}\}_{0\leq t\leq 1}$ be a path in $\mathrm{SL}(2,\mathbb{R})$ such that $\phi_{t}(c_{j})=U^{j}_{t}\phi(c_{j})(U^{j}_{t})^{-1}.$ We extend $\phi_{t}$ from $\pi_{1}(M)$ to $\pi_{1}(\Sigma)$ by defining $\phi_{t}(\gamma)=U^{j}_{t}\phi(\gamma)(U^{j}_{t})^{-1}$ for all $\gamma\in\pi_{1}(N_{j})$. This defines a continuous path $\{\phi_{t}\}_{0\leq t\leq 1}$ of representations in $\mathcal{R}^{\prime}_{h}(\Sigma)$ such that $\phi_{1}(c)$ is hyperbolic.
Now the restriction of $\phi_{1}$ to $\pi_{1}\left(\Sigma^{\prime}\right)$ lies in $\mathcal{R}^{\prime}_{h}(\Sigma^{\prime})$, where $\Sigma^{\prime}=N_{1}\cup N_{2}\cup M_{2}$ is a subsurface of $\Sigma$. By the induction hypothesis, there exists a path in $\mathcal{R}_{h}^{\prime}(\Sigma^{\prime})$ from this representation to one which maps each interior boundary in decomposition of $\Sigma^{\prime}$ to a hyperbolic element. Furthermore, by the path-lifting property (see Lemma 4.9), this path can be extended to a path in $\mathcal{R}_{h}^{\prime}(\Sigma)$.
Similarly, if there is a component $N_{3}$ attached to the boundary $B$, then, as in the cases of $N_{1}$ and $N_{2}$, we can also show that $\phi$ can be deformed into an interior hyperbolic representation in $N_{3}$.
This completes the proof. ∎
Lemma 4.11.
Let $\Sigma=\Sigma_{g,n}$ be a surface of genus $g\geq 1$ with $n\geq 1$ boundary components, and let $\Sigma=\bigcup_{i=1}^{|\chi(\Sigma)|}M_{i}$ be a maximal dual tree decomposition of $\Sigma$. Then any representation $\phi$ such that each $\phi|_{\pi_{1}(M_{i})}$ is non-abelian can be deformed to an interior hyperbolic representation $\phi^{\prime}$ while keeping the boundary conjugacy classes fixed. Moreover, the deformations is non-abelian on each $M_{i}$.
Proof.
The argument is similar to the proof of Lemma 4.10. Since $g\geq 1$, we choose $M_{1}$ to be a punctured torus, and $M_{2}$ to be a pair of pants.
By Lemma 4.7, the representation $\phi$ can be deformed to $\phi_{1}$ such that $\phi_{1}(c)$ is hyperbolic. Moreover, $\phi_{1}$ remains non-abelian when restricted to each $M_{i}$.
Now the restriction of $\phi_{1}$ to $\pi_{1}\big{(}\bigcup_{i>1}M_{i}\big{)}$ lies in $\mathcal{R}^{\prime}_{h}(\Sigma^{\prime})$, where $\Sigma^{\prime}=\bigcup_{i>1}M_{i}$ is a subsurface of $\Sigma$. From Lemma 4.10, $\phi_{1}|_{\pi_{1}(\Sigma^{\prime})}$ can be deformed to an interior hyperbolic representation. By the path–lifting property [10, Corollary 7.8], these deformations can be extended to a path in $\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SL}(2,\mathbb{R}))$. This completes the proof. ∎
4.2.2. The maximal Toledo invariant
@@
Let $\Sigma_{g,n}$ be a surface of genus $g\geq 1$ with $n$ boundary components, and let $\mathrm{Par}(\Sigma_{g,n})$ denote the space of boundary parabolic representations. Then
@@
We first consider the case where $n\geq 2$. Since $g\geq 1$, by Lemma 4.11, we may assume that $C_{0},D_{1},\dots,D_{n-2}$ are all hyperbolic.
@@
$\{\mathrm{T}^{-1}(|\chi(\Sigma_{g,n})|)\cap\mathrm{Par}(\Sigma_{g,n})\}=2^{2g+n-1}$
for $g\geq 1$.
@@
4.3. The case of $n$ is even and $g\geq 1$
Let $\Sigma_{g,n}$ be a surface of genus $g\geq 1$ with $n$ boundary components. Let $\mathrm{Hyp}(\Sigma_{g,n})$ denote the space of boundary hyperbolic representations in $\mathrm{Hom}(\pi_{1}(\Sigma_{g,n}),\mathrm{SL}(2,\mathbb{R}))$. For a hyperbolic element $C\in\mathrm{SL}(2,\mathbb{R})$, recall the sigma map is $\sigma(C)=\frac{1}{2}(1+\mathrm{sgn}(\mathrm{tr}(C)))$.
@@ 4.3. The case of $n$ is even and $g\geq 1$
(4.20) $\{\widehat{\mathcal{P}}_{g,n}\}=8^{n/2}\cdot\{\#\widehat{\mathcal{H}}_{g,n/2}\}=8^{n/2}\cdot 2^{n/2-1}(4g+n-3)=2^{2n-1}(4g+n-3).$
@@
Let $\phi\in\mathrm{Hom}(\pi_{1}(\Sigma_{g,n}),\mathrm{SL}(2,\mathbb{R}))$ be a representation such that $\phi(c_{1}),\dots,\phi(c_{n-1})$ are parabolic and $\phi(c_{n})$ is hyperbolic, $g\geq 1$. Then there exists a deformation $\phi(t)$ of $\phi$ such that the conjugacy classes of $\phi(t)(c_{1}),\dots,\phi(t)(c_{n-1})$ remain fixed, the trace $\mathrm{tr}(\phi(t)(c_{n}))$ varies monotonically, and $\phi(1)(c_{n})$ is parabolic with $\mathrm{tr}(\phi(1)(c_{n}))=2\mathrm{sgn}(\mathrm{tr}(\phi(c_{n})))$.
@@
We first assume $n\geq 2$. By a small perturbation fixing the conjugacy classes of $\phi(c_{1}),\dots,\phi(c_{n-1})$, we may assume that $\phi(\Pi_{i=1}^{g}[a_{i},b_{i}]c_{1}\cdots c_{n-2})$ is either elliptic or hyperbolic. This can be done since $g\geq 1$. Up to conjugation, we may write
@@
(4.22) $\mathrm{T}(\phi)=\mathrm{T}(\psi)+\tfrac{1}{2}s_{a},$
@@
Let $\phi\in\mathrm{Par}(\Sigma_{g,n})$ with even $n\geq 2$ and $g\geq 1$. Then $\phi$ and $\gamma\phi$ lie in the same connected component of $\mathrm{Par}(\Sigma_{g,n})$ for any $\gamma\in\Lambda_{g,n}$ if and only if $\mathrm{T}(\psi)$ is neither maximal nor minimal.
@@
For any $\phi_{0},\phi_{1}\in\{\phi\}$, the proof of Lemma 4.16 shows that the corresponding representations $\psi_{0}$ and $\psi_{1}\in\mathrm{Hyp}(\Sigma_{g,n/2})$ have the same Toledo invariant and the same sigma values.
@@
If $\mathrm{T}(\psi_{0})=\mathrm{T}(\psi_{1})$ is maximal or minimal, then $\psi_{0}$ and $\gamma\cdot\psi_{1}$ are connected for some $\gamma\in\Lambda_{g,n}$ by the fact that there are $4^{g}$ different components with the same sigma value for $\mathrm{Hyp}(\Sigma_{g,n/2})$ coming from the action of $\Lambda_{g,n}$. From the proof of Lemma 4.20, we know that each of the elements $\phi^{\prime}_{i}(a_{j})$ and $\phi^{\prime}_{i}(b_{j})$, for $i=1,2$ and $1\leq j\leq g$, cannot be elliptic. Since $\phi^{\prime}_{0}$ and $\phi^{\prime}_{1}$ are connected, the corresponding elements $\psi(a_{j})=\phi^{\prime}_{i}(a_{j})$ and $\psi(b_{j})=\phi^{\prime}_{i}(b_{j})$ must lie in the closure of the same hyperbolic region. This implies that $\gamma=\mathrm{id}$, and hence $\psi_{0}$ and $\psi_{1}$ are connected.
@@
Hence, the number of connected components of the space $\mathrm{Par}(\Sigma_{g,n})$ of boundary-parabolic representations, assuming $n\geq 2$ is even and $g\geq 1$, is given by
@@ 4.4.1. Connected components of $\mathrm{HP}(\Sigma_{g,n})$
For any $a=(a_{1},a_{2},a_{3})\in\{0,1\}\times\{0,1\}\times\{\pm 1,\pm\sqrt{-1}\}$, the subset $\sigma^{-1}(a)\subset\mathrm{HP}(\Sigma_{0,3})$ is non-empty and connected. In particular, the space $\mathrm{HP}(\Sigma_{0,3})$ has $16$ connected components.
@@
Suppose $c\neq 0$. If $a_{1}+a_{2}+\mathrm{Re}(a_{3})$ is even, then $\mathrm{tr}(C_{2}C_{3})=(-1)^{a_{1}}(a+d+cs_{3})$. One can choose parameters $a,d,c$ such that $a+d>2$ and $a+d+cs_{3}<-2$. Hence $C_{1}=(C_{2}C_{3})^{-1}$ is hyperbolic and satisfies $\sigma(C_{1})=a_{1}$. In particular, in this case, we have
(4.25) $cs_{3}<-(a+d)-2<0,$
which implies that the sign of $c$ is uniquely determined by $s_{3}$.
Similarly, when $a_{1}+a_{2}+\mathrm{Re}(a_{3})$ is odd, we can choose $a,d,c$ such that $a+d+cs_{3}>2$ and again $C_{1}$ is hyperbolic with the correct sigma value. In both cases, we conclude that $\sigma^{-1}(a)$ is non-empty.
@@
Now suppose that $\sum_{i=1}^{3}\mathrm{Re}(a_{i})$ is even. Let $\phi,\phi^{\prime}\in\sigma^{-1}(a)$ be two representations with corresponding parameters $c$ and $c^{\prime}$. By (4.34), we have $cc^{\prime}>0$. Similarly to the case when $\sum_{i=1}^{3}\mathrm{Re}(a_{i})$ is odd, we can show that $\phi$ and $\phi^{\prime}$ lie in the same connected component. This completes the proof. ∎
@@
[…]
$\mathrm{T}(\phi)=\begin{cases}1&\text{for }a=(0,1,-1),\ (1,0,-1),\ (0,0,\sqrt{-1}),\ (1,1,\sqrt{-1}),\\ -1&\text{for }a=(0,1,1),\ (1,0,1),\ (0,0,-\sqrt{-1}),\ (1,1,-\sqrt{-1}),\\ 0&\text{for }\sum_{i=1}^{3}\mathrm{Re}(a_{i})\ \text{odd}.\end{cases}$
@@
In both cases, the conjugacy classes of $\phi(c_{i})$, $1\leq i\leq n-2$, are unchanged during the deformation, $\phi(c_{n-1})$ is perturbed slightly, and $\phi(c_{n})$ becomes hyperbolic with sigma value $|\mathrm{Re}(\sigma(\phi(c_{n})))|$ for all $t>0$. Hence, the Toledo invariant remains constant throughout the deformation.
@@
is well-defined. Moreover, it is a double covering when restricted to $\mathcal{HP}_{g,n}\cap\mathrm{T}^{-1}(k)$ with $|k|<|\chi(\Sigma_{g,n})|$, and is of degree one when restricted to $\mathcal{HP}_{g,n}\cap\mathrm{T}^{-1}(\pm|\chi(\Sigma_{g,n})|)$.
@@
We now prove that $\Psi$ is a covering map when restricted to $\mathrm{T}^{-1}(k)$ and determine its degree. Consider a pair of pants $(C_{1},C_{2},C_{3})$ with $C_{1}$, $C_{2}$ and $C_{3}$ hyperbolic. By (3.3), there is a deformation of $C_{3}$ to a parabolic element, keeping the conjugacy class of $C_{1}$ and the sigma values of $C_{2}$ fixed. During this deformation, the trace of $C_{3}$ maintains its sign, and the Toledo invariant remains constant. In particular, by Lemma 3.2, if $\sum_{i=1}^{3}\mathrm{Re}\big{(}\sigma(C_{i})\big{)}$ is odd (i.e., the Toledo invariant is zero), then the conjugacy class of the parabolic element can be $\pm\sqrt{\mathrm{sgn}(\mathrm{tr}(C_{3}))}$. If $\sum_{i=1}^{3}\mathrm{Re}\big{(}\sigma(C_{i})\big{)}$ is even (i.e., the Toledo invariant is $\pm 1$), then the conjugacy class of the parabolic element is unique (see the proof of Lemma 4.23 or Remark 4.24).
For any $\phi^{\prime}\in\mathrm{Hyp}(\Sigma_{g,n})$ with $n\geq 2$ and $g\geq 1$, if $|\mathrm{T}(\phi^{\prime})|<|\chi(\Sigma_{g,n})|$, then by Lemma 3.5 we can deform $\phi^{\prime}$ to a new representation $\phi^{\prime\prime}\in\mathrm{Hyp}(\Sigma_{g,n})$ which is still interior hyperbolic and satisfies that $\phi^{\prime\prime}|_{\pi_{1}(P_{n})}$ has Toledo invariant zero, where $P_{n}$ is a pair of pants containing $c_{n}$. From the above discussion for pairs of pants, there exist at least two representations $\phi_{0},\phi_{1}\in\mathrm{HP}(\Sigma_{g,n})$ with different $\sigma$–values on $c_{n}$ such that $\Psi(\{\phi_{0}\})=\Psi(\{\phi_{1}\})=\{\phi^{\prime\prime}\}=\{\phi^{\prime}\}.$ If $|\mathrm{T}(\phi^{\prime})|=\pm|\chi(\Sigma_{g,n})|$, then, by the above discussion, there exists a unique representation $\phi_{0}\in\mathrm{HP}(\Sigma_{g,n})$ such that $\Psi(\{\phi_{0}\})=\{\phi^{\prime}\}.$
Suppose $\phi_{0},\phi_{1}\in\mathrm{HP}(\Sigma_{g,n})\cap\sigma^{-1}(a)\cap\Psi^{-1}(\{\phi^{\prime}\})$, then $\phi_{0}^{\prime}$ and $\phi_{1}^{\prime}$ are connected by a path $\phi^{\prime}(t)\in\mathrm{Hyp}(\Sigma_{g,n})$. Reversing the deformation from $\phi^{\prime}(t)(c_{n})$ back to a parabolic element (as in Lemma 4.25), we obtain a path $\phi(t)$ in $\mathrm{HP}(\Sigma_{g,n})\cap\sigma^{-1}(a)$ connecting $\phi_{0}$ and $\phi_{1}$. Thus, the fiber $\Psi^{-1}(\{\phi^{\prime}\})\cap\sigma^{-1}(a)$ is connected, and so has at most one connected component. If $|\mathrm{T}(\phi^{\prime})|<|\chi(\Sigma_{g,n})|$, then the $\sigma$–values on $c_{n}$ distinguish $\phi_{0}$ and $\phi_{1}$, and the entire fiber consists of exactly two elements. If $|\mathrm{T}(\phi^{\prime})|=|\chi(\Sigma_{g,n})|$, then the $\sigma$–values of $\phi_{0}$ and $\phi_{1}$ coincide, and the entire fiber consists of exactly one element.
The argument for the case $n=1$ is analogous: the deformation occurs on the boundary curve $c_{1}$, while the part corresponding to $\Pi_{i=1}^{g-1}[a_{i},b_{i}]$ remains unchanged, and only a slight perturbation is performed on $a_{g},b_{g}]$, which remains hyperbolic throughout the deformation. This allows the construction to proceed in the same way. ∎
@@
(4.26) $\begin{split}&\quad\#\{\mathrm{HP}(\Sigma_{g,n})\}\\ &=2\cdot\#\{\mathrm{Hyp}(\Sigma_{g,n})\cap\mathrm{T}^{-1}(k),|k|<|\chi(\Sigma_{g,n})|\}\\ &\quad+\{\mathrm{Hyp}(\Sigma_{g,n})\cap\mathrm{T}^{-1}(\pm|\chi(\Sigma)|)\}\\ &=2^{n}(4g+2n-5)+4^{g}\cdot 2^{n}\\ &=2^{n}(4^{g}+4g+2n-5).\end{split}$
for any $(g,n)\neq(1,1)$ and $g\geq 1$.